a fence forms a rectangle and uses 64 meters of fencing.if the width of the enclosed area is 13 meters what is the length
Answer by macston(5194)

Perimeter=2(L+W)
64 m=2(L+13m) Divide each side by 2
32m=L+13m Subtract 13m from each side.
19m=L ANSWER: The length is 19 meters.